October 20, 2022
RC Meeting Information
RC Daily Meetings are at 11:00 (am) via Zoom and in person in the counting house meeting room (2nd floor)
- Zoom remote connection information:
Join ZoomGov Meeting https://jlab-org.zoomgov.com/j/1607100885?pwd=ZFBsRTFsYm5Zb0laZ0tyZXNjcEE0QT09 Meeting ID: 160 710 0885 Passcode: 595071 One tap mobile +16692545252,,1607100885# US (San Jose) +16468287666,,1607100885# US (New York) Dial by your location +1 669 254 5252 US (San Jose) +1 646 828 7666 US (New York) +1 669 216 1590 US (San Jose) +1 551 285 1373 US 833 568 8864 US Toll-free Meeting ID: 160 710 0885 Find your local number: https://jlab-org.zoomgov.com/u/ab2CekLXgQ
- SHMS: 8 degrees, -9.2 GeV, 1x, 2x 1-7 off
- HMS: 20 degrees, -6.6 GeV
5pass beam, 60uA when possible. Hall C has beam priority ;)
- Continue data taking at setting 1
- beam trip cut?
- adjust statistical goals?
- date for ladder change?
Work in the hall
- SHMS preshower channel 13 moved to a new HV channel: https://logbooks.jlab.org/entry/4063287
- target loop 2 fan readback fixed: https://logbooks.jlab.org/entry/4062317
- SHMS aerogel turned off: https://logbooks.jlab.org/entry/4063605
- Jixie and Xinzhan replaced a gas bottle for the SoLID GEM test. They will need make another entrance next week.
- Around 5:30pm, BCM05 (UNSER) temperature started to fluctuated (from 110 to 130 degree) https://logbooks.jlab.org/entry/4063493, MCC called EES expert to come in. The control * * card in counting house was replaced (https://logbooks.jlab.org/entry/4063553) and UNSER went back to the set temperature at 11pm (https://logbooks.jlab.org/entry/4063610)
- a new UNSER calibration?
- BCM calibration
- Install Q1 SHMS vacuum gauge (~3hr)